# Bilayer Excitons in the Laughlin Fractional Quantum Hall State

### v2

@@
Notably, the interlayer excitons observed along the Jain sequence should be distinguished from those described by the (111) state at $\nu_{1}=\nu_{2}=1/2$ Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Liu2022crossover , as well as those associated with two-component fractional quantum Hall effect Zhang2025fractionalexciton . Ground states described by two-component wavefunctions, such as the (111) and (333) states Halperin1983 ; Wen1992 , arise in regimes where strong interlayer correlation plays a critical role in defining the FQHE phase. In contrast, the regime explored in this work is characterized by weak interlayer coupling, where FQHE states are stabilized primarily by intralayer Coulomb interactions.
The regimes of strong and weak interlayer correlation can be identified based on the hierarchical behavior of the FQHE sequence Jain.03 ; Li2019pairing . The boundary between these two regimes are revealed in Figure 4e, which plots $G_{\text{PF}}$ as a function of LL filling and $B$, measured along the equal density line. The regime of strong interlayer correlation occurs at $B<10$ T, where the FQHE sequence aligns with composite fermions incorporating interlayer flux attachment Li2019pairing ; Zhang2025fractionalexciton . In contrast, at $B>15$ T, the FQHE sequence exhibits the characteristic hierarchical structure of the Jain sequence. These two regimes are separated by a series of $B$-induced transitions at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, marked by black horizontal lines in Fig. 4e, where the charge gap is weakened due to a transition between two-component and one-component FQHE ground states.
Amongst the Jain sequence of FQHE states in the regime of weak interlayer correlation, transport measurements unveil an unexpected FQHE state at $\nu_{1}=\nu_{2}=3/8$. This state emerges at $B=12$ T and persists beyond $B=31$ T (Fig. 4b-e), offering strong support that the underlying charge order is single-component in nature. However, $\nu_{1}=\nu_{2}=3/8$ lies outside the Jain sequence of FQHE, indicating an exotic type of order.
In the following, we propose a plausible description of relevant physics at $\nu_{1}=\nu_{2}=3/8$.
We first discuss the ground state wavefunction in each graphene layer, which defines the nature of the charge gap. Since this filling corresponds to half-filled $\Lambda$ level of component fermions  Pan.03 ; xia2004electron ; pan2015fractional ; bellani2010optical ; samkharadze2015observation , the emergence of a robust FQHE state could follow a similar mechanism as the celebrated $\nu=5/2$ state. In this case, electrons in the lowest LL capture two-flux quanta and populate $\Lambda$ levels in response to the residual magnetic field. When the second $\Lambda$ level is half-filled, the two-flux composite fermions capture two additional flux quanta, undergo Cooper instability, and condensed into a $p-$wave paired state. This scenario is supported by previous numerical simulations, identifying a relative to the Bonderson-Slingerland (BS) wavefunction as the stable ground state at $\nu=3/8$ in the lowest LL of a single 2D confinement  scarola2002possible ; Mukherjee2012Pfaffian ; mukherjee2014possible ; Bonderson2008 . This describes a non-Abelian state similar to the Pfaffian state at $\nu=5/2$  Halperin2020fractional . In this scenario, quasiparticle excitations in the sample bulk likely obeys non-Abelian quantum statistics, which arises from the non-trivial fusion rules between the distinct topological sectors.
Most interestingly, the transport response at $\nu_{1}=\nu_{2}=3/8$ offers clear indication for interlayer exciton pairing. Besides vanishing $G_{\text{PF}}$, this state exhibits an enhanced counterflow conductance, as well as a peak in the drag ratio approaching unity. Given the robust charge gap and the single-component nature of the ground state, we argue that exciton pairing occurs between quasiparticle and quasihole excitations. Notably, if the FQHE state at $\nu_{1}=\nu_{2}=3/8$ is non-Abelian in nature, the resulting exciton will follow non-Abelian statistics as well, giving rise to a highly unusual construction (see Methods section for more detailed discussion on the likely non-Abelian order of the bilayer state).
Previous studies on ultra-high mobility GaAs samples have reported the FQHE at $\nu=3/8$, originating from residual composite fermion interaction  wang2022even ; wang2023next ; wang2023fractional . However, this exotic state has been notably absent in prior experimental studies involving monolayer graphene samples with similar doubly-encapsulated geometry  Zibrov2018even ; Zeng2019 ; Chen2019 ; Polshyn2018corbino . It should be noted that LL mixing plays a significant role in facilitating interaction between composite fermions, creating favorable condition for pairing at half-filled $\Lambda$ levels  zhao2023composite ; wang2023next . Indeed, these unconventional states in GaAs were observed in hole-doped LL filling rather than electron-doped, presumably because the larger effective mass of holes, and consequently larger LL mixing, enhances interactions between CFs wang2022even ; wang2023next ; wang2023fractional . In quantum Hall graphene bilayers, we expect that the unique Coulomb environment reduces short-range repulsion in the CF sea, making it more favorable for CFs to form Cooper pairs. At large magnetic field, the absence of interlayer correlations restricts Cooper pairing to occur within each individual layer. As a result, the bilayer system offers a favorable platform to stabilize novel composite fermion states driven by intra-layer correlations, which are distinct from those observed in monolayer graphene samples.
In summary, we report a new type of bilayer exciton in quantum Hall graphene bilayers, constructed through pairing between quasiparticles and quasiholes of single-component FQHE states. At the Laughlin state, fractionally charged quasiparticles exhibit a braiding phase of $32\pi/$ 3. Because the particle and hole components of the exciton reside in separate layers, the resulting exciton inherits a total braiding phase of $34\pi/$ 3. This leads to a highly unusual composite particle that is overall charge-neutral but obeys anyonic quantum statistics. The flow dynamics of neutral anyons can be directly explored using counterflow transport, while their population can be uniquely controlled through layer-asymmetric field-effect doping. Together, this new type of bilayer exciton open a new frontier for engineering exotic quantum phases, poised to inspire broad future research directions.
@@ acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledges partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.
@@
N.J.Z. and J.I.A.L. conceived the project. N.J.Z., R.Q.N., and X.L. fabricated the device. N.J.Z. and R.Q.N. performed the measurement. N.B. and D.E.F. provided theoretical inputs. K.W. and T.T. provided the material. N.J.Z., R.Q.N., N.B., D.E.F., and J.I.A.L wrote the manuscript together.

### v3

@@
Notably, the interlayer excitons observed along the Jain sequence should be distinguished from those described by the $(111)$ state at $\nu_{1}=\nu_{2}=1/2$ Eisenstein2014; Li2017superfluid; Liu2017superfluid; Liu2022crossover, as well as from those associated with two-component fractional quantum Hall states Zhang2025fractionalexciton. Ground states characterized by two-component wavefunctions, such as the $(111)$ and $(333)$ states Halperin1983; Wen1992, emerge in regimes where strong interlayer correlations play a defining role in establishing the FQH phase. In contrast, the Laughlin state is stabilized predominantly by intralayer Coulomb interactions, while excitons in this regime arise from weak interlayer coupling between quasiparticle excitations.
Exciton pairing at $3/8+3/8$ filling
Having established exciton pairing in the Laughlin state, we now extend our investigation to other FQH states along the equal-density line defined by $\nu_{1}=\nu_{2}=\nu$. Fig. 4a presents the parallel-flow conductance $G_{\mathrm{PF}}$ as a function of $\nu$ and $B$ along the equal-density line, revealing a transition between regimes of strong and weak interlayer correlation (see Fig. M4 for a full range map). FQH states with strong interlayer coupling appear at $B<10$ T, where the positions of insulating states show excellent agreement with two-component FQH states with interlayer flux attachment Li2019pairing; Zhang2025fractionalexciton. As $B$ increases, a crossover from two-component to single-component FQH states leads to a suppression in the charge gap, producing a series of $B$-induced transitions at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, as indicated by the horizontal black lines in Fig. 4a. In the high-field regime of weak interlayer correlation, the FQH states follow the characteristic hierarchical structure of the Jain series, consistent with wavefunctions that are single-component in nature.
Interestingly, in the regime of weak interlayer correlation, a FQH state emerges at $\nu_{1}=\nu_{2}=3/8$, which lies outside the conventional Jain sequence Pan.03; xia2004electron; pan2015fractional; bellani2010optical; samkharadze2015observation. This state not only exhibits clear signatures of exciton pairing but also displays a drag response that is qualitatively distinct from that of the surrounding Jain fractions (Fig. 4b–d). As illustrated in Fig. 4c, the state at $\nu_{1}=\nu_{2}=3/8$ shows a substantially enhanced counterflow conductance compared with nearby fillings. Moreover, along the equal-density line, the $\nu_{1}=\nu_{2}=3/8$ state produces the most prominent counterflow drag signal (Fig. 4e), underscoring the unusually strong interlayer correlations at this filling. Strikingly, the drag signal at $\nu_{1}=\nu_{2}=3/8$ is sharply confined to $\Delta\nu=0$: as shown in Fig. 4f, the drag response decays rapidly once $\Delta\nu$ departs from zero, in stark contrast to the broader response observed at $\nu_{1}=\nu_{2}=1/3$ (Fig. 2b–g).
Exciton pairing at $\nu_{1}=\nu_{2}=3/8$ represents an unprecedented phenomenon. While the drag response by itself does not allow an unambiguous identification of the underlying ground state, we discuss below two plausible scenarios.
First, exciton pairing may be an intrinsic component of the underlying FQH state. Canonical examples include the $(111)$ state at $\nu_{1}=\nu_{2}=1/2$ and the $(333)$ state at $\nu_{1}=\nu_{2}=1/6$ Wen1992; Eisenstein2014; Zhang2025fractionalexciton. Along this vein, the simplest construction for $\nu_{1}=\nu_{2}=3/8$ is an effective $(111)$-like state formed between half-filled $\Lambda$-levels (see Methods for further discussion). The precise structure of such a state, however, remains unclear and will require future theoretical work.
Alternatively, the exciton may share a similar construct to the Laughlin state at $\nu_{1}=\nu_{2}=1/3$. In this scenario, the FQH state is primarily stabilized by intra-layer Coulomb interactions, while exciton pairing arises from binding a quasiparticle in one graphene layer to a quasihole in the other. Since this filling corresponds to a half-filled $\Lambda$-level of composite fermions Pan.03; xia2004electron; pan2015fractional; bellani2010optical; samkharadze2015observation, the FQH state is likely described by a non-Abelian wavefunction  scarola2002possible; Mukherjee2012Pfaffian; mukherjee2014possible; Bonderson2008.
In the second scenario, quasiparticle excitations are non-abelian, yielding excitons that inherit non-Abelian statistics. The unusually strong interlayer correlations at this filling may therefore be a manifestation of non-Ableian topological order. Notably, although a FQH state at $3/8$ filling has recently been reported in ultra–high-mobility GaAs samples wang2022even; wang2023next; wang2023fractional, it has been conspicuously absent in monolayer graphene Zibrov2018even; Zeng2019; Chen2019; Polshyn2018corbino. As such, further theoretical and experimental investigations will be required to resolve the nature of this state.
Taken together, our findings reveal a new direction in which exciton pairing serves as the pathway for realizing novel quantum particles that behave like charge-neutral anyons. This new class of bilayer excitons opens a promising frontier for engineering exotic quantum phases, including anyonic superfluidity Han2025anyon, and is poised to stimulate broad future research efforts.
@@ acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledge partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. S.A. acknowledges support from Kuwait University. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.
@@
N.J.Z. and J.I.A.L. conceived the project. N.J.Z., R.Q.N., and X.L. fabricated the device. N.J.Z., R.Q.N., S.A. performed the measurement. N.B. and D.E.F. provided theoretical inputs. K.W. and T.T. provided the material. N.J.Z., R.Q.N., N.B., S.A., D.E.F., and J.I.A.L wrote the manuscript together.